OpenStudy (anonymous):

Can someone please check my answers? Will metal and fan. Which decimal is equivalent to 7.44343434… ? my answer > 7.4434

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

would you mind checking a few more?

OpenStudy (anonymous):

thank you by the way. :)

OpenStudy (mathmath333):

y w

OpenStudy (anonymous):

Which is not a factor pair of 400? 6 and 55

OpenStudy (mathmath333):

none

OpenStudy (anonymous):

huh ?

OpenStudy (mathmath333):

400=2^4*5^2

OpenStudy (mathmath333):

6=2*3 55=5*11

OpenStudy (anonymous):

yeah i get that. the options are 6 and 55 8 and 50 16 and 25 20 and 20 my answer was 6 and 55

OpenStudy (mathmath333):

6*55=330 8*50=400 16*25=400 20*20=400